Lindy sold her Fender acoustic guitar on eBay using i-soldit.com. The i-soldit location she used charges 35% of the first $500 of the selling price and 20% of the amount over $500. After these charges were deducted from the selling price, she received $745. For how much did her guitar sell?
Answer by rwm(914)

let s= selling price
o= amt of selling price over 500
v =fee on over 500
500*.35=175
s-500=o
.2*o=v
s-175-v=745
o = 525, s = 1025, v = 105
selling price 1025
check
1025-175-105=745
ok